### v1

@@ Abstract
We performed large, lab-in-the-field experiment (2,591 participants across 134 Honduran villages; ten rounds) and tracked how contribution behavior unfolds in fixed, anonymous groups of size five. Contribution separates early into two durable paths, one low and one high, with rare convergence thereafter. High-path players can be identified with strong accuracy early on. Groups that begin with an early majority of above‑norm contributors (about 60%) are very likely finish high. The empirical finding of a bifurcation, consistent with the theory, shows that early, high contributions by socially central people steer groups onto, and help keep them on, a high-cooperation path.
@@ 1 Introduction
Why do some communities sustain costly cooperation while others succumb to chronic free-riding ostrom1990governing? Local public goods (like clean water systems, school maintenance, road repairs, and vaccination drives) depend on repeated voluntary contributions that are individually costly and thus hard to explain with standard self-interest. Laboratory public goods games (PGGs) typically show a decline in average contributions over time. Such experiments (e.g. henrich2010markets) also reveal striking and persistent heterogeneity in behavior: some individuals never contribute, some contribute consistently, and many settle into intermediate patterns. This variation often appears trait, like across contexts, in a “cooperative phenotype” peysakhovich2014humans. Here, we examine whether persistent differences in cooperation across otherwise similar rural villages reflect individuals’ and villages’ characteristics.
@@
Lab PGGs typically involve low stakes (e.g., fehr2000cooperation), and simply raising stakes does not, on average, increase cooperation (kocher2008does). In high-salience settings, however, larger stakes combined with informational or comparative feedback can boost contributions, see liu2024signaling (China), cardenas2015between (Colombia), and nhim2023water (Cambodia). In this spirit, our high-salience field setting, substantial stakes relative to local incomes, plus repeated feedback, provides a demanding test of how peer exposure and individual traits shape contribution dynamics. Our experimental design and panel data structure allow us to identify the effect of peers and individual traits on contribution dynamics. In particular, we exploit the panel nature of the game with ten rounds of observations and use specifications with individual fixed effects and village-by-round fixed effects. By incorporating leave-one-out lagged peer contribution averages, we can separate a player’s responsiveness to peers from their own persistent cooperativeness, mitigating concerns about reflection or unobserved heterogeneity in group composition. This approach leverages within-person changes over time and across-round exposure to peer behavior to isolate the drivers of sustained cooperation.
@@
We have three primary findings. First, across village, contribution trajectories bifurcate into two distinct and durable paths, high versus low cooperation, rather than converging to a single interior equilibrium. A small early advantage in a person’s peer environment (i.e., being surrounded by others who contribute in the initial rounds) is enough to push that individual onto a high-contribution path; lacking that, behavior tends to settle into persistently low contributions. At the village level, the probability of “finishing high” (i.e. sustaining a high-cooperation group by the final round) rises sharply once the early share of high contributors exceeds a threshold, with a data-driven critical mass around 60% of players contributing above average in the first few rounds. This pattern is indicative of strategic complementarities or positive feedback: early cooperation begets more cooperation, locking the group into a high-contribution equilibrium, whereas an early deficit of cooperators leads to a low-contribution trap.
@@
Second, network attributes, particularly the number of friendship ties measures prior to the game, but not adversarial ties, strongly predict entry into and persistence on the high-cooperation path, such that each additional friend is associated with higher average contributions and individuals who are more embedded (and less enmeshed in conflicts) are more likely to sustain cooperation. Personal attributes also matter: men contribute about 0.6 Lempiras more per round than women on average, and more schooling is positively associated with giving (modestly higher average contributions across schooling levels). Cultural identity plays a role as well; relative to Catholics (the majority religion in our sample), non-religious villagers and Protestants contribute about 0.4 and 0.2 additional Lempiras per round, respectively. Finally, age by itself has at most a limited effect on cooperation once we account for social ties.
@@
These patterns are not driven by any single threshold choice or by terminal-round outcomes. Dynamics are path-dependent, downward moves from high to low (drop $H\!\to\!L$) are more common than rebounds from low to high (rise $L\!\to\!H$). Friendship ties buffer against $H\!\to\!L$ drops, and reporting no religion is associated with $L\!\to\!H$ rises. Round-to-round updating places greater weight on others’ prior contributions than on one’s own past choice; higher education weakens simple repetition and increases adjustment toward the group, and men place more weight on peers. These qualitative patterns are robust across specifications, and a variance decomposition indicates that most residual variation lies at the individual level, with little remaining structure at the village level after accounting for observed covariates.
@@ 2 Setting and Experimental Design
Before play, we mapped within-village social networks using standard name-generator questions airoldi2024induction. Each respondent identified (i) a set of friends (e.g. people they spend time with, ask advice of, or would borrow/lend small amounts of money to) and (ii) adversarial ties (people they avoid, have conflicts with, or “dislike”). These lists produced, at the individual level, friend degree and adversarial degree, and, at the village level, summary measures (e.g., average friend degree). We measured education as years of schooling completed. Catholics constitute the majority of villages, with observed minorities of Protestants and non-religious respondents. We also collected other standard individual-level measures, including age, gender, marital status, education, indigenous status, and food insecurity (as a measure of player wealth). Village-level measures involved remoteness (access routes), village size, friendship density, and adversarial density.
Sessions consisted of ten identical rounds of a linear public-goods game (PGG) played in fixed groups of five participants drawn from the same village. In each round, every player received an endowment of up to 12 Lempiras and chose how much to contribute to a common pool. The sum of contributions was doubled and then evenly redistributed among group members. Choices were recorded privately; at the end of the session, participants were paid the sum of their round-by-round earnings. Participants were not told the total number of rounds ex ante to avoid horizon effects and within-each-round choices were simultaneous and private; feedback on group outcomes arrived only after the round closed.
@@ 4.1 Stage Game and Instantaneous Payoff
  • Norm aversion: $h_{i}\exp\!\bigl[k(c_{it}-\bar{c}_{-i,t-1})^{2}\bigr]$, where $h_{i}\in[0,1]$ measures the salience of conforming to the lagged peer norm $\bar{c}_{-i,t-1}:=\frac{1}{N-1}\sum_{j\neq i}c_{j,t-1}$, and $k>0$ controls curvature. [^5] Any smooth, bounded penalty that is minimized at zero deviation yields qualitatively identical results.
@@
$\pi_{i}^{t}=\frac{b}{N}\Bigl[c_{it}+(N-1)\bar{c}_{-i,t}\Bigr]-\kappa\,c_{it}+d_{i}\bigl(c_{it}\bigr)^{\alpha}-h_{i}\exp\!\Bigl[-k\bigl(c_{it}-\bar{c}_{-i,t-1}\bigr)^{2}\Bigr].$ (1)
@@ 4.2 Best-Reply Behavior
$D(c)=\Bigl(\tfrac{b}{N}-\kappa\Bigr)+d_{i}\,\alpha\,c^{\alpha-1}+2k\,h_{i}\,(c-\bar{c}_{-i,t-1})\exp\!\bigl[-k(c-\bar{c}_{-i,t-1})^{2}\bigr]=0.$
Because $\alpha<1$, the altruism term is strictly concave up deceasing; the norm aversion term is monotone in $h_{i}$ and pulls the optimum toward $\bar{c}_{-i,t-1}$. The best reply is single-valued, continuous, and bounded.
@@ 4.3 Adaptive Dynamics
Solving $D(c)=0$ we obtain the singular strategy $c^{*}$. The stability of $c^{*}$ is determined by the $D^{\prime}(c^{*})$. $c^{*}$ is convergence stable if $D^{\prime}(c^{*})<0$.
@@ 4.4 Convergence Stability and Evolutionary Branching
$d_{i}\,\alpha(\alpha-1)\,(c^{*})^{\alpha-2}+2k\,h_{i}<0.$
@@
The population initially converges to the singular strategy $c^{*}$ by following the selection gradient. However, since $c^{*}$ is a local fitness minimum (not satisfying the Evolutionarily Stable Strategy condition), ”mutants” can invade $c^{*}$ both from below and above, leading the onset of evolutionary branching. The population splits into two paths, one toward high contribution versus the other toward low contribution.
@@ 4.6 Identification and Empirical Mapping
Only the product $kh_{i}$ enters the ESS second-order condition, whereas both $d_{i}$ and $h_{i}$ affect the first-order condition. Separate identification of $(d_{i},h_{i},k)$ therefore leverages auxiliary measures and the panel:
  • $h_{i}$, we conceptually interpret it as norm‑aversion; in our empirical mapping we calibrate $kh_{i}$ from dispersion.
  • $k$ is calibrated from the cross-round dispersion of $(c_{it}-\bar{c}_{-i,t-1})^{2}$.
  • $d_{i}$ is estimated by maximum likelihood given $(h_{i},k)$, with variance-inflation checks.
@@ 4.7 Instrumental–variables strategy
For the same–round peer mean $\bar{c}_{-i,gt}$ we use leave–one–out (LOO) composition instruments constructed from predetermined traits $W^{(k)}$:
$Z_{ig}^{(k)}\;=\;\frac{1}{N-1}\sum_{j\neq i}W_{j}^{(k)}.$
@@
To capture learning, we target the lagged peer mean $\bar{c}_{-i,g,t-1}$ and use a single time–varying shift–share instrument. Let $s^{(k)}_{g,-i}$ be the LOO group share of trait $k$, and let $\mu^{(k)}_{t-1}$ be the round–(t1)(t\!-\!1) cross–village mean of that trait. We implement a leave–one–village (LOV) variant that excludes village $v(i)$ from the shifter:
$Z^{LOV}_{igt}\;=\;\sum_{k}s^{(k)}_{g,-i}\,\mu^{(-v),(k)}_{t-1}$, (5)
@@ 5.2 Heterogeneity
Each additional named friend is associated to slightly higher giving with +0.03+0.03 L ($p=0.039$). This pattern is consistent with a natural social capital mechanism: having more friends may heighten reciprocity pressures, increase reputational visibility, and expose individuals to stronger pro-social norms. Although the estimated effect is small for any single tie, it can accumulate for highly connected players. Because the specification is in levels, however, the coefficient likely blends both “preference” and “exposure” channels, and may also proxy for centrality or social status.
@@
We cluster complete ten‑round histories using Ward–D2 on the zz‑scored player multiplied by round matrix. We restrict to players with all ten observed rounds and zz‑score contributions by round before clustering to avoid scale differences across rounds. Average silhouette width peaks at two clusters ($\bar{S}=0.362$) and falls thereafter ($\bar{S}=0.258,0.218,0.176,0.152$ for $k=3,\dots,6$; Table 2). At $k=2$ we obtain two cohesive groups (Table 3): a larger High cluster (n=1,826n=1{,}826, $\bar{S}=0.34$) and a smaller Low cluster ($n=765$, $\bar{S}=0.42$). Alignment with round‑10 end states is strong (Table 4): the High cluster contains 1,041 of 1,055 HH finishers (98.7%) and most movers (330/472 HL; 179/195 LH), whereas the Low cluster is predominantly LL (593/848=69.9%) and rarely HH (14/848=1.7%). Cluster mean paths remain well separated across rounds (Figure 4).
@@
Table 3: Cluster sizes and mean silhouette widths at $k=2$
@@
Figure 4: Mean zz‑scored contribution paths by Ward cluster ($k=2$). The High path starts above the round‑wise mean and remains high; the Low path starts below and declines.
@@ Appendix A Theory & Proofs
$Z^{(k)}_{ig}$ LOO composition IV: mean of trait $k$ among $i$’s groupmates.
$s^{(k)}_{g,-i}$ LOO share of trait $k$ in $i$’s group.
$\mu^{(k)}_{t-1}$ Cross-village round-(t1)(t\!-\!1) mean (shifter) for trait $k$.
$Z^{SS}_{igt}$ Shift–share IV: $\sum_{k}s^{(k)}_{g,-i}\,\mu^{(k)}_{t-1}$.
@@
$\pi_{i}^{t}=\underbrace{\frac{b}{N}\big(c_{i,t}+(N-1)\bar{c}_{-i,t}\big)-\kappa\,c_{i,t}}_{\text{material payoff}}\;+\;\underbrace{d_{i}\,(c_{i,t})^{\alpha}}_{\text{altruism}}\;-\;\underbrace{h_{i}\,\exp\!\big[-k\,(c_{i,t}-\bar{c}_{-i,t-1})^{2}\big]}_{\text{norm aversion}}\$
@@
$D(c)\;=\;\left(\frac{b}{N}-\kappa\right)\;+\;d_{i}\,\alpha\,c^{\alpha-1}\;+\;2k\,h_{i}\,(c-\bar{c}_{-i,t-1})\,\exp\!\big[-k(c-\bar{c}_{-i,t-1})^{2}\big]\;=\;0.$
@@
where the norm-aversion term vanishes at $c_{m}=c$ because $\partial/\partial c_{m}\big[-h_{i}\exp\{-k(c_{m}-c)^{2}\}\big]=+2kh_{i}(c_{m}-c)\exp\{-k(c_{m}-c)^{2}\}=0$ at $c_{m}=c$. Then $D$ has a unique zero $c^{*}\in(0,12]$ given by
@@
$\left.\frac{\partial^{2}\pi_{i}^{t}(c_{m};\,c)}{\partial c_{m}^{2}}\right|_{c_{m}=c=c^{*}}\;=\;d_{i}\,\alpha(\alpha-1)\,(c^{*})^{\alpha-2}\;+\;2k\,h_{i}.$
@@
$\pi_{i}^{t}(c_{m};\,c)=\Big(\tfrac{b}{N}-\kappa\Big)c_{m}\;+\;\tfrac{b(N-1)}{N}\,c\;+\;d_{i}\,c_{m}^{\alpha}\;-\;h_{i}\exp\!\big\{-k(c_{m}-c)^{2}\big\}.$
@@
$\frac{\partial\pi}{\partial c_{m}}=\Big(\tfrac{b}{N}-\kappa\Big)\;+\;d_{i}\,\alpha\,c_{m}^{\alpha-1}\;+\;2k\,h_{i}(c_{m}-c)\,\exp\!\big\{-k(c_{m}-c)^{2}\big\}.$
@@
$\frac{\partial^{2}\pi}{\partial c_{m}^{2}}=d_{i}\,\alpha(\alpha-1)\,c_{m}^{\alpha-2}\;+\;2k\,h_{i}\cdot\Big[\tfrac{\mathrm{d}}{\mathrm{d}c_{m}}\big((c_{m}-c)\,e^{-k(c_{m}-c)^{2}}\big)\Big].$
The bracket equals $1$ at $c_{m}=c$ (since $\frac{\mathrm{d}}{\mathrm{d}x}[xe^{kx^{2}}]\big|_{x=0}=1$), hence
$\left.\frac{\partial^{2}\pi}{\partial c_{m}^{2}}\right|_{c_{m}=c}=d_{i}\,\alpha(\alpha-1)\,c^{\alpha-2}+2k\,h_{i}.$

### v2

@@ Abstract
We performed large lab-in-the-field experiment (2,591 participants across 134 Honduran villages; ten rounds) and tracked how contribution behavior unfolds in fixed, anonymous groups of size five. Contributions separates early into two durable paths, one low and one high, with rare convergence thereafter. High-path players can be identified with strong accuracy early on. Groups that begin with an early majority of above‑norm contributors (about 60%) are very likely finish to high. The empirical finding of a bifurcation, consistent with the theory, shows that early, high contributions by socially central people steer groups onto, and help keep them on, a high-cooperation path.
@@ 1 Introduction
Why do some communities sustain costly cooperation while others succumb to chronic free-riding ostrom1990governing? Local public goods (like clean water systems, school maintenance, road repairs, and vaccination drives) depend on repeated voluntary contributions that are individually costly and thus hard to explain with standard self-interest. Laboratory public goods games (PGGs) typically show a decline in average contributions over time. Such experiments (e.g. henrich2010markets) also reveal striking and persistent heterogeneity in behavior: some individuals never contribute, some contribute consistently, and many settle into intermediate patterns. This variation often appears trait-like across contexts, in a “cooperative phenotype” peysakhovich2014humans. Here, we examine whether persistent differences in cooperation across otherwise similar rural villages reflect individuals’ and villages’ characteristics.
@@
Lab PGGs typically involve low stakes (e.g., fehr2000cooperation), and simply raising stakes does not, on average, increase cooperation (kocher2008does). In high-salience settings, however, larger stakes combined with informational or comparative feedback can boost contributions, see liu2024signaling (China), cardenas2015between (Colombia), and nhim2023water (Cambodia). In this spirit, our high-salience field setting, substantial stakes relative to local incomes, plus repeated feedback, provide a demanding test of how peer exposure and individual traits shape contribution dynamics. Our experimental design and panel data structure allow us to identify the effect of peers and individual traits on contribution dynamics. In particular, we exploit the panel nature of the game with ten rounds of observations and use specifications with individual fixed effects and village-by-round fixed effects. By incorporating leave-one-out lagged peer contribution averages, we can separate a player’s responsiveness to peers from their own persistent cooperativeness, mitigating concerns about reflection or unobserved heterogeneity in group composition. This approach leverages within-person changes over time and across-round exposure to peer behavior to isolate the drivers of sustained cooperation.
@@
We have three primary findings. First, across villages, contribution trajectories bifurcate into two distinct and durable paths, high versus low cooperation, rather than converging to a single interior equilibrium. A small early advantage in a person’s peer environment (i.e., being surrounded by others who contribute in the initial rounds) is enough to push that individual onto a high-contribution path; lacking that, behavior tends to settle into persistently low contributions. At the village level, the probability of “finishing high” (i.e. sustaining a high-cooperation group by the final round) rises sharply once the early share of high contributors exceeds a threshold, with a data-driven critical mass around 60% of players contributing above average in the first few rounds. This pattern is indicative of strategic complementarities or positive feedback: early cooperation begets more cooperation, locking the group into a high-contribution equilibrium, whereas an early deficit of cooperators leads to a low-contribution trap.
@@
Second, network attributes, particularly the number of friendship ties measured prior to the game, but not adversarial ties, strongly predict entry into and persistence on the high-cooperation path, such that each additional friend is associated with higher average contributions and individuals who are more embedded (and less enmeshed in conflicts) are more likely to sustain cooperation. Personal attributes also matter: men contribute about 0.6 Lempiras more per round than women on average, and more schooling is positively associated with giving (modestly higher average contributions across schooling levels). Cultural identity plays a role as well; relative to Catholics (the majority religion in our sample), non-religious villagers and Protestants contribute about 0.4 and 0.2 additional Lempiras per round, respectively. Finally, age by itself has at most a limited effect on cooperation once we account for social ties.
@@
These patterns are not driven by any single threshold choice or by terminal-round outcomes. Dynamics are path-dependent: downward moves from high to low (drop $H\!\to\!L$) are more common than rebounds from low to high (rise $L\!\to\!H$). Friendship ties buffer against $H\!\to\!L$ drops, and reporting no religion is associated with $L\!\to\!H$ rises. Round-to-round updating places greater weight on others’ prior contributions than on one’s own past choice; higher education weakens simple repetition and increases adjustment toward the group, and men place more weight on peers. These qualitative patterns are robust across specifications, and a variance decomposition indicates that most residual variation lies at the individual level, with little remaining structure at the village level after accounting for observed covariates.
@@ 2 Setting and Experimental Design
Before play, we mapped within-village social networks using standard name-generator questions (see airoldi2024induction). Each respondent identified (i) a set of friends (e.g. people they spend time with, ask advice of, or would borrow/lend small amounts of money to) and (ii) adversarial ties (people they avoid, have conflicts with, or “dislike”). These lists produced, at the individual level, friend degree and adversarial degree, and, at the village level, summary measures (e.g., average friend degree). We measured education as years of schooling completed. Catholics constitute the majority of villages, with observed minorities of Protestants and non-religious respondents. We also collected other standard individual-level measures, including age, gender, marital status, education, indigenous status, and food insecurity (as a measure of player wealth). Village-level measures involved remoteness (access routes), village size, friendship density, and adversarial density.
Sessions consisted of ten identical rounds of a linear public-goods game (PGG) played in fixed groups of five participants drawn from the same village. In each round, every player received an endowment of 12 Lempiras and chose how much to contribute to a common pool. The sum of contributions was doubled and then evenly redistributed among group members. Choices were recorded privately; at the end of the session, participants were paid the sum of their round-by-round earnings. Participants were not told the total number of rounds ex ante to avoid horizon effects and within each round choices were simultaneous and private; feedback on group outcomes arrived only after the round closed.
@@ 4.1 Stage Game and Instantaneous Payoff
  • Aversion to being at the norm: $h_{i}\exp\!\bigl[-k_{\text{norm}}(c_{it}-\bar{c}_{-i,t-1})^{2}\bigr]$, and $h_{i}\in[0,1]$ measures the salience of anti-conforming to the lagged peer norm $\bar{c}_{-i,t-1}:=\frac{1}{N-1}\sum_{j\neq i}c_{j,t-1}$, and $k_{\text{norm}}>0$ controls curvature. [^5] Any smooth, bounded cost function $P(c_{it}-\bar{c}_{-i,t-1})$ that attains its maximum at zero deviation (so the utility term $P(\cdot)$ is minimal there) yields qualitatively similar results. [^6] Any smooth, bounded penalty that is maximized at zero deviation yields qualitatively identical results.
@@
$\pi_{i}^{t}=\frac{b}{N}\Bigl[c_{it}+(N-1)\bar{c}_{-i,t}\Bigr]-\kappa\,c_{it}+d_{i}\bigl(c_{it}\bigr)^{\alpha}-h_{i}\exp\!\Bigl[-k_{\text{norm}}\bigl(c_{it}-\bar{c}_{-i,t-1}\bigr)^{2}\Bigr].$ (1)
@@ 4.2 Best-Reply Behavior
$D(c)=\Bigl(\tfrac{b}{N}-\kappa\Bigr)+d_{i}\,\alpha\,c^{\alpha-1}+2k_{\text{norm}}\,h_{i}\,(c-\bar{c}_{-i,t-1})\exp\!\bigl[-k_{\text{norm}}(c-\bar{c}_{-i,t-1})^{2}\bigr]=0.$
Because $\alpha<1$, the altruism term is strictly decreasing and concave up in $c$; the aversion to being at the norm term is monotone in $h_{i}$ and pulls the optimum away $\bar{c}_{-i,t-1}$. For the parameter range we consider, the best reply is numerically single-valued, continuous, and bounded on $[0,12]$.
@@ 4.3 Adaptive Dynamics
Solving $D(c)=0$ we obtain the singular strategy $c^{*}$. The stability of $c^{*}$ is determined by $D^{\prime}(c^{*})$. $c^{*}$ is convergence stable if $D^{\prime}(c^{*})<0$.
@@ 4.4 Convergence Stability and Evolutionary Branching
$d_{i}\,\alpha(\alpha-1)\,(c^{*})^{\alpha-2}+2k_{\text{norm}}\,h_{i}<0.$
@@
The population initially converges to the singular strategy $c^{*}$ by following the selection gradient. However, when the ESS condition fails, that is, when
$d_{i}\,\alpha(\alpha-1)(c^{*})^{\alpha-2}+2k_{\text{norm}}h_{i}>0,$
$c^{*}$ becomes a local fitness minimum, so “mutants” close to $c^{*}$ on both sides can invade, leading to the onset of evolutionary branching.
@@ 4.6 Identification and Empirical Mapping
Only the product $\phi_{i}=2k_{\text{norm}}h_{i}$ enters the ESS second-order (ESS/branching) condition, whereas the first-order condition for the singular strategy $c^{*}$ depends only on $d_{i}$. Separate identification of $(d_{i},\phi_{i})$ therefore leverages auxiliary measures and the panel:
  • $h_{i}$, we conceptually interpret it as aversion to being at the norm; in our empirical mapping we calibrate $kh_{i}$ from dispersion.
  • $k_{\text{norm}}$ is calibrated from the cross-round dispersion of $(c_{it}-\bar{c}_{-i,t-1})^{2}$.
  • $d_{i}$ is estimated by maximum likelihood given $(h_{i},k_{\text{norm}})$, with variance-inflation checks.
@@ 4.7 Instrumental–variables strategy
For the same–round peer mean $\bar{c}_{-i,gt}$ we use leave–one–out (LOO) composition instruments constructed from predetermined traits $W^{(l)}$:
$Z_{ig}^{(l)}\;=\;\frac{1}{N-1}\sum_{j\neq i}W_{j}^{(l)}.$
@@
To capture learning, we target the lagged peer mean $\bar{c}_{-i,g,t-1}$ and use a single time–varying shift–share instrument. Let $s^{(l)}_{g,-i}$ be the LOO group share of trait $l$, and let $\mu^{(l)}_{t-1}$ be the round–(t1)(t\!-\!1) cross–village mean of that trait. We implement a leave–one–village (LOV) variant that excludes village $v(i)$ from the shifter:
$Z^{LOV}_{igt}\;=\;\sum_{l}s^{(l)}_{g,-i}\,\mu^{(-v),(l)}_{t-1}$, (5)
@@ 5.2 Heterogeneity
Each additional named friend is associated with slightly higher giving with +0.03+0.03 L ($p=0.039$). This pattern is consistent with a natural social capital mechanism: having more friends may heighten reciprocity pressures, increase reputational visibility, and expose individuals to stronger pro-social norms. Although the estimated effect is small for any single tie, it can accumulate for highly connected players. Because the specification is in levels, however, the coefficient likely blends both “preference” and “exposure” channels, and may also proxy for centrality or social status.
@@
We cluster complete ten‑round histories using Ward–D2 on the zz‑scored player multiplied by round matrix. We restrict to players with all ten observed rounds and zz‑score contributions by round before clustering to avoid scale differences across rounds. Average silhouette width peaks at two clusters ($\bar{S}=0.362$) and falls thereafter ($\bar{S}=0.258,0.218,0.176,0.152$ for $k=3,\dots,6$; Table 2). At two clusters ($K=2$) we obtain two cohesive groups (Table 3): a larger High cluster (n=1,826n=1{,}826, $\bar{S}=0.34$) and a smaller Low cluster ($n=765$, $\bar{S}=0.42$). Alignment with round‑10 end states is strong (Table 4): the High cluster contains 1,041 of 1,055 HH finishers (98.7%) and most movers (330/472 HL; 179/195 LH), whereas the Low cluster is predominantly LL (593/848=69.9%) and rarely HH (14/848=1.7%). Cluster mean paths remain well separated across rounds (Figure 4).
@@
Table 3: Cluster sizes and mean silhouette widths at two clusters (KK=2)
@@
Figure 4: Mean zz‑scored contribution paths by Ward cluster (two clusters $K=2$). The High path starts above the round‑wise mean and remains high; the Low path starts below and declines.
@@ Appendix A Theory & Proofs
$Z^{(l)}_{ig}$ LOO composition IV: mean of trait $l$ among $i$’s groupmates.
$s^{(l)}_{g,-i}$ LOO share of trait $l$ in $i$’s group.
$\mu^{(l)}_{t-1}$ Cross-village round-(t1)(t\!-\!1) mean (shifter) for trait $l$.
$Z^{SS}_{igt}$ Shift–share IV: $\sum_{l}s^{(l)}_{g,-i}\,\mu^{(l)}_{t-1}$.
@@
$\pi_{i}^{t}=\underbrace{\frac{b}{N}\big(c_{i,t}+(N-1)\bar{c}_{-i,t}\big)-\kappa\,c_{i,t}}_{\text{material payoff}}\;+\;\underbrace{d_{i}\,(c_{i,t})^{\alpha}}_{\text{altruism}}\;-\;\underbrace{h_{i}\,\exp\!\big[-k_{\text{norm}}\,(c_{i,t}-\bar{c}_{-i,t-1})^{2}\big]}_{\text{aversion to being at the norm}}\$
@@
$D(c)\;=\;\left(\frac{b}{N}-\kappa\right)\;+\;d_{i}\,\alpha\,c^{\alpha-1}\;+\;2k_{\text{norm}}\,h_{i}\,(c-\bar{c}_{-i,t-1})\,\exp\!\big[-k_{\text{norm}}(c-\bar{c}_{-i,t-1})^{2}\big]\;=\;0.$
@@
where the aversion to being at the norm term vanishes at $c_{m}=c$ because $\partial/\partial c_{m}\big[-h_{i}\exp\{-k_{\text{norm}}(c_{m}-c)^{2}\}\big]=+2k_{\text{norm}}h_{i}(c_{m}-c)\exp\{-k_{\text{norm}}(c_{m}-c)^{2}\}=0$ at $c_{m}=c$. Then $D$ has a unique zero $c^{*}\in(0,12]$ given by
@@
$\left.\frac{\partial^{2}\pi_{i}^{t}(c_{m};\,c)}{\partial c_{m}^{2}}\right|_{c_{m}=c=c^{*}}\;=\;d_{i}\,\alpha(\alpha-1)\,(c^{*})^{\alpha-2}\;+\;2k_{\text{norm}}\,h_{i}.$
@@
$\pi_{i}^{t}(c_{m};\,c)=\Big(\tfrac{b}{N}-\kappa\Big)c_{m}\;+\;\tfrac{b(N-1)}{N}\,c\;+\;d_{i}\,c_{m}^{\alpha}\;-\;h_{i}\exp\!\big\{-k_{\text{norm}}(c_{m}-c)^{2}\big\}.$
@@
$\frac{\partial\pi}{\partial c_{m}}=\Big(\tfrac{b}{N}-\kappa\Big)\;+\;d_{i}\,\alpha\,c_{m}^{\alpha-1}\;+\;2k_{\text{norm}}\,h_{i}(c_{m}-c)\,\exp\!\big\{-k_{\text{norm}}(c_{m}-c)^{2}\big\}.$
@@
$\frac{\partial^{2}\pi}{\partial c_{m}^{2}}=d_{i}\,\alpha(\alpha-1)\,c_{m}^{\alpha-2}\;+\;2k_{\text{norm}}\,h_{i}\cdot\Big[\tfrac{\mathrm{d}}{\mathrm{d}c_{m}}\big((c_{m}-c)\,e^{-k_{\text{norm}}(c_{m}-c)^{2}}\big)\Big].$
The bracket equals $1$ at $c_{m}=c$ (since $\frac{\mathrm{d}}{\mathrm{d}x}[xe^{k_{\text{norm}}x^{2}}]\big|_{x=0}=1$), hence
$\left.\frac{\partial^{2}\pi}{\partial c_{m}^{2}}\right|_{c_{m}=c}=d_{i}\,\alpha(\alpha-1)\,c^{\alpha-2}+2k_{\text{norm}}\,h_{i}.$